OpenStudy (anonymous):

Can anyone pls help me with this? Solve the logarithmic equation algebraically. Approximate the result to three decimal places. (If there is no solution, enter NO SOLUTION.) ln x + ln(x + 7) = 3

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

Try combining the natural logs, then do something to both sides of the equation

OpenStudy (anonymous):

In(x(x+7)) = 3

OpenStudy (anonymous):

and e^in (x(x+7)) = e^3, right?

zepdrix (zepdrix):

yes very good c:

OpenStudy (anonymous):

then x^2 + 7x = e^3?

OpenStudy (anonymous):

thank you for the medal, @math>philosophy

zepdrix (zepdrix):

Yes, good. And since we have this ugly exponential term, this won't end up factoring nicely. We'll have to throw it into the quadractic formula :D

OpenStudy (anonymous):

how?

zepdrix (zepdrix):

x^2 + 7x - e^3 = 0 This is of the form: ax^2 + bx + c = 0 Remember how to solve for x? :D

OpenStudy (anonymous):

yes i do but what to do w/ the e?

OpenStudy (anonymous):

Remember e is just a constant, like pi

zepdrix (zepdrix):

You certainly could punch it into your calculator if you wanted to e^3 = 20.08553692... Or you can leave it in as an exponential til the end. Your answer might turn out more accurately if you don't use the decimal right now :D

OpenStudy (anonymous):

ok i will solve it and let u know

OpenStudy (anonymous):

is it 2.186?

zepdrix (zepdrix):

Hmm I came up with different numbers. Lemme see what went wrong.

zepdrix (zepdrix):

\[\large 0=x^2+7x-e^3\] \[\large x=\frac{-7 \pm \sqrt{49+4e^3}}{2}\]

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

I got ~2.186 as one of the solutions

OpenStudy (anonymous):

yea i got that too

zepdrix (zepdrix):

Ok i musta made a mistake :) If you have that same setup for your quadratic formula, then you're probably doing it correctly.

OpenStudy (anonymous):

its ok :)

zepdrix (zepdrix):

2.186 yah that sounds right. And of course, you should have 2 solutions, since it's plus/minus the sqrt. But that value doesn't lie in the domain of our original function since we can't take the natural log of a negative value! :O You didn't mention that answer, so maybe you already figured that out :D

OpenStudy (anonymous):

i got it right, thnx to u an @math>philosophy